How do I solve this problem? x(3x + 1) (x + 1) = 0
Answer by Fombitz(32388)

x(3x + 1) (x + 1) = 0
You need to find x so that the equation equals zero.
There are three cases when the equation will equal zero, that is, when the three individual expressions equal zero.
1.x=0
2.3x+1=0
3.x+1=0
The first solution is x=0.
The second solution requires some manipulating.
3x+1=0
3x=-1
x=-1/3.
The second solution is x=-1/3.
The third solution is found similarly.
x+1=0
x=-1
The third solution is x=-1.